OpenStudy (anonymous):

does anyone have the awnsers for Lesson 1: Rate of Change and Slope Algebra 1 A Unit 5: Linear Functions

OpenStudy (anonymous):

i bet i know them name a question

OpenStudy (anonymous):

i love algebra.....shoot the q

OpenStudy (anonymous):

A line has slope –5/3. Through which two points could this line pass? (1 point)(12, 13), (17, 10) (16, 15), (13, 10) (0, 7), (3, 10) (11, 13), (8, 18)

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

i havnt the slightest clue i dont even know how to do it

OpenStudy (anonymous):

you need to compute four slopes for this one problem

OpenStudy (anonymous):

and how do u do that

OpenStudy (anonymous):

but we can cheat a little i think in general the slope is \(\frac{y_2-y_1}{x_2-x_1}\) now the first one the numerator would be \(10-13=-3\) so forget that one

OpenStudy (hba):

@omgineedhlp I guess i answered the same ques in your last post.

OpenStudy (anonymous):

for \[(16, 15), (13, 10)\] we compute \[\frac{10-15}{13-16}=\frac{-5}{-3}=\frac{5}{3}\] so it is not that one

OpenStudy (anonymous):

lets try \[(11, 13), (8, 18)\] \[\frac{18-13}{8-11}=\frac{5}{-3}=-\frac{5}{3}\] got it on the last try

OpenStudy (anonymous):

ok so 11,13 the 11 is x1?

OpenStudy (anonymous):

The pair of points (6, y) and (10, –1) lie on a line with slope . What is the value of y? (1 point)–5 –2 2 5